### v1

# Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models

Abstract
Text-to-Image (T2I) models often suffer from issues such as semantic leakage, incorrect feature binding, and omissions of key concepts in the generated image. This work studies these phenomena by looking into the role of information flow between textual token representations. To this end, we generate images by applying the diffusion component on a subset of contextual token representations in a given prompt and observe several interesting phenomena. First, in many cases, a word or multiword expression is fully represented by one or two tokens, while other tokens are redundant. For example, in “San Francisco’s Golden Gate Bridge”, the token “gate” alone captures the full expression. We demonstrate the redundancy of these tokens by removing them after textual encoding and generating an image from the resulting representation. Surprisingly, we find that this process not only maintains image generation performance, but also reduces errors by 21% compared to standard generation. We then show that information can also flow between different expressions in a sentence, which often leads to semantic leakage. Based on this observation, we propose a simple, training-free method to mitigate semantic leakage: replacing the leaked item’s representation after the textual encoding with its uncontextualized representation. Remarkably, this simple approach reduces semantic leakage by 85%. Overall, our work provides a comprehensive analysis of information flow across textual tokens in T2I models, offering both novel insights and practical benefits. [^1] We release our code at https://github.com/tokeron/lens
1 Introduction
Text-to-image (T2I) models consist of two main components: a text encoder and a diffusion model (Ho et al., 2020; Song & Ermon, 2019). The former processes the user’s prompt, transforming it into a representation that guides the latter in generating the image. While T2I models are powerful and widely used, misalignment issues frequently arise. Generated images often fail to capture key concepts from the user’s prompt, leading to catastrophic neglect (Chefer et al., 2023a), semantic leakage of properties between entities (Rassin et al., 2022), or incorrectly bind different concepts (Huang et al., 2023; Rassin et al., 2022). Prior work has attempted to address these issues by modifying the cross-attention mechanism between the diffusion model and the encoded textual tokens (Rassin et al., 2023; Chefer et al., 2023a; Dahary et al., 2024), under the implicit assumption that the text encoder accurately captures the user’s intent and that each textual token primarily represents the item it is intended to convey.
In this work, we challenge these assumptions and investigate how information flows within tokens during the textual encoding process, and what information is encoded in each textual token representation. [^2] Throughout this work, we study token representations at the output of the text encoder. For brevity, we sometimes omit the word “representation”. To this end, we analyze lexical items within prompts, where each lexical item is a single word or multiword expression that functions as a unit of meaning. Our analysis focuses on two complementary aspects. First, in-item information flow—how different sub-word token representations of the same lexical item encode and distribute information. For example, do all the token representations of the item “pelican” (“pe”, “lic”, and “an”) encode this lexical item (Fig. 1, top left). Second, inter-item information flow—how information flows between different lexical items (Feng et al., 2024; Feng & Steinhardt, 2024). For example, given the prompt “A businesswoman on a runway”, does the representation of “runway” change in the context of the word “businesswoman” (Fig. 1, bottom middle).
Figure 1: Our main findings. Left: Information within a lexical item is unevenly distributed across its token representations, with a few carrying most of the meaning. Middle: Context may cause incorrect information flow between items, distorting their representations and leading to unintended images. In this example, the representation of “runway” is distorted in the context of “a businesswoman”. Right: We propose a simple solution—encode the leaked item (“runway” in this case) separately and patch it into the representation of the prompt after textual encoding.
For in-item information flow (4), we adapt a causal intervention method (Toker et al., 2025; see Section 3.2) to generate images from individual token representations. Specifically, we first encode the entire sentence, and then generate an image from each individual contextualized token representation. The resulting image reflects the information that the diffusion model can reconstruct when conditioned on that single token’s contextual representation. Our experiments with FLUX (black-forest labs, 2024) show that, in most cases, each lexical item is fully represented by only one or two tokens, while its remaining tokens might be considered redundant (4.2). Finally, we demonstrate that our findings can be used to improve image–text alignment in image generation. To this end, we mask all the non-representative tokens (tokens that do not represent the lexical item) just before the diffusion, and generate images directly from the remaining representations. We find that this operation alone reduces image generation error by 21% compared to standard image generation.
For inter-item information flow (5), we adapt the same causal intervention method—this time to assess the information encoded in one lexical item with respect to others in the prompt. For instance, in the sentence a “A pool and a table”, we examine the information flow between the items “pool” and “table”. We find that in 89% of cases, each lexical item’s information is preserved and not influenced by others. However, in some cases, information flows between tokens. Interestingly, this flow can occur between lexical items that should not influence each other, for example, “table” influencing the representation of “pool” in the sentence above, changing pool’s representation to a “pool table”. We show that this leakage can be mitigated by patching the leaked lexical item’s representation with an uncontextualized representation of this item. Notably, this simple intervention prevents leakage in 85% of cases.
In summary, we analyze information flow in T2I models and uncover two key findings: (1) lexical items are often represented by just one or two tokens, and removing non-representative tokens can improve generation; (2) while different lexical items usually do not influence each other’s representations, when they do, it can happen between items that should not influence each other—leading to semantic leakage. Finally, based on our observation, we propose a simple method to mitigate semantic leakage, leading to a 85% reduction in leakage errors.
2 Related work
2.1 Interpretability
Interpretability in multimodal models.
Previous work on interpretability in T2I models analyzed the text encoder. Chefer et al. (2023b) examined the concepts encoded in the latent space of CLIP (Radford et al., 2021) and how different concepts relate to each other. Toker et al. (2024) studied the progression of textual representations through the layers of the text encoder, using the diffusion model as a lens into the encoding process. Another line of work explored the connection between the text encoder and the image generator (Tang et al., 2023). In this work we investigate how information flows between tokens within a single layer. We focus on the final layer, as it directly conditions the diffusion process.
Information flow in LLMs.
Interpretability research has examined how token information propagates via attention in LLMs (Vig & Belinkov, 2019; Clark et al., 2019). However, raw attention weights can be misleading (Pruthi et al., 2020; Jain & Wallace, 2019). To address this, methods such as Attention Rollout (Abnar & Zuidema, 2020) were developed to more reliably measure information flow through attention. Other studies have employed Patchscopes (Ghandeharioun et al., 2024) and the logit lens (nostalgebraist, 2020) to interpret the encoding processes in LLMs. These works demonstrate that modern language models go far beyond simple subword processing. For instance, Kaplan et al. (2025) showed that models perform an implicit detokenization process, effectively fusing subword tokens into coherent word-level representations. Feucht et al. (2024) reported an “erasure” effect, in which the contributions of earlier tokens are overwritten by later ones—suggesting the emergence of an internal lexicon. Another prominent direction is probing (Adi et al., 2016; Liu et al., 2019). For example, Zhang & Bowman (2018); Brunner et al. (2019) trained probes to classify, from a token’s intermediate representation, the identity of tokens a fixed number of positions away. However, these methods are not fully reliable, as probes can exploit spurious correlations (Belinkov, 2022). Moreover, these works typically assess whether information is present or flows between tokens, but not whether the next step in the pipeline (e.g., the LM head or decoder) actually uses this information. In our analysis, we focus on intervention methods that test whether the next component in the image generation pipeline can retrieve and use the relevant information.
2.2 Challenges and solutions in multimodal models.
The recent rise of T2I models has introduced several challenges, prompting various approaches to address them. This work shows that a better understanding of information flow between tokens can help mitigate many of these challenges.
Semantic leakage.
One notable challenge in T2I models is semantic leakage (Rassin et al., 2022). For example, given the prompt “a bat flying in a baseball stadium”, the generated image might include a baseball bat instead of a bat (the animal). A similar effect has been observed in language models, where the context influences generation even when it conceptually should not (Gonen et al., 2024).
Attribute binding.
Another issue is attribute binding, which concerns the correct association of attributes with entities mentioned in the prompt. For example, in the prompt “a yellow flamingo and a pink sunflower”, models must correctly bind “yellow” to “flamingo” and “pink” to “sunflower”. Several approaches have been proposed to address these challenges, including performing an optimization in the latent space to produce an attention pattern that is better aligned with the syntactic structure of the sentence (Rassin et al., 2023), and merging tokens to concentrate information into a single representation (Hu et al., 2024).
Catastrophic neglect.
A further prominent challenge is catastrophic neglect, where entities mentioned in the prompt are completely omitted from the generated image. Many works have attempted to mitigate this, one of the first being Attend and Excite (Chefer et al., 2023a), which optimizes the image latents to increase attention on the neglected entity tokens.
3 Methodology
3.1 Intervention on the text encoder
Our goal in this paper is to evaluate the effect of information flow between textual token representations, which later condition the diffusion process. We are interested in analyzing how information is distributed across tokens within a lexical item (e.g., do all the tokens in “San Francisco’s Golden Gate Bridge” represent that item, or is the information more diffusely spread?). To do so, we must be able to interpret the information encoded in individual tokens. We are also interested in measuring how different lexical items influence one another. To this end, we seek to isolate a subset of tokens and evaluate their joint representation. For these purposes, we adjust text intervention in diffusion models (Toker et al., 2025), allowing us to generate images from arbitrary subsets of tokens within the encoded user prompt.
Given a prompt with $N$ tokens, $t_{1},t_{2},\ldots,t_{N}$, our goal is to isolate and interpret the information encoded by a subset of these tokens. Let $S\subset\{1,\ldots,N\}$ be the index set of selected tokens, where $1\leq|S|\leq N$. We begin by encoding the full prompt using the text encoder $E$, yielding the final hidden states $e_{1},\ldots,e_{N}$. Separately, we encode a sequence consisting entirely of pad tokens to obtain pad embeddings $p_{1},\ldots,p_{N}$. We then construct a patched prompt by replacing all hidden states outside $S$ with the corresponding pad embeddings:
$\tilde{t}_{i}=\begin{cases}e_{i}&\text{if }i\in S,\\ p_{i}&\text{otherwise}\end{cases}\quad\text{for }i=1,\ldots,N.$
The patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ is then used to guide the diffusion model. Generating an image from this patched representation allows us to visualize and isolate the individual contributions of the selected tokens as interpreted by the diffusion model. To evaluate the information in the generated images, we use a vision-language model (VLM): For the inter-item information flow experiments, we assess whether the generated image represents the full lexical item, while for the in-item information flow experiments, we measure whether information from other lexical items is present in the generated image. See Section 3.2 for an illustration of the method applied to interpret individual subtoken representations.
3.2 Experimental setup
Models.
We experiment with FLUX-Schnell (black-forest labs, 2024)—a recent state-of-the-art T2I model. We also report similar findings on FLUX-dev in A.5.1. FLUX models employ T5-XXL (Raffel et al., 2019) as their text encoder, enabling bidirectional information flow between tokens. In A.5.2 we also include another model, SDXL-Turbo (Sauer et al., 2023), which uses a different text encoder—CLIP (Radford et al., 2021). Unlike T5-xxl, CLIP is a causal language model. See Appendix A.5.2 for the analysis of SDXL-Turbo.
Data.
We use a subset of prompts from DrawBench (Saharia et al., 2022) and Parti (Yu et al., 2022b). In total, we use 1,053 prompts. For each prompt, we generate five images using different random seeds. To conduct the analysis in our paper, we first find the lexical items in each prompt. We use GPT-4o (OpenAI et al., 2024) to extract the lexical items, resulting in a total of 4,864 unique lexical items. See A.1 for the exact prompt we used and further technical details. We then use spaCy (Honnibal et al., 2020) to determine the part of speech of each lexical item, and retain only nouns, proper nouns, and adjectives, as these are typically concrete and can be identified in their visual representation. We end up with 3,891 unique lexical items and use them to evaluate both inter-item information flow and in-item information flow.
Evaluation.
To evaluate the content of generated images, we also employ Qwen2-VL-72B-Instruct (Wang et al., 2024), a model with strong general vision capabilities. We restrict our evaluation to binary (yes/no) questions. See A.1 for the exact prompt used and additional details.
Figure 2: –ֿ
Evaluating in-item information flow: Our proposed framework interprets the information flow within a lexical item. We generate images from each token comprising the lexical item (left-hand side) and analyze the resulting images using a VLM (right-hand side). In this example, we interpret the different tokens composing the item pelican and find that only one of them, lic, represents the concept of a pelican, while the other two tokens (“pe”, and “an”) do not.
4 In-item information flow
4.1 How is information distributed across tokens?
In this section, we explore how information is distributed across the token representations within a lexical item. We aim to answer two main questions: Do all tokens encode the same information? And is the information evenly distributed across tokens or concentrated in specific token representations? These questions are central, as many T2I applications (Chefer et al., 2023a; Rassin et al., 2023; Dahary et al., 2024) treat all tokens of a lexical item equally. Discovering a non-uniform distribution of information across tokens could improve application effectiveness by focusing on the most informative tokens. We apply our intervention method (Section 3.1) to the hidden representation of each token in the prompt. We then generate an image conditioned on each token’s contextualized representation. We use Qwen2-VL to assess whether an image represents the lexical item it is part of. We repeat this analysis for each lexical item, for each prompt in our combined dataset (Section 3.2). The results show that in 89% of the cases, at least one of the tokens fully represents the lexical item. Out of the remaining 11%, about 85% cannot be generated even from the full lexical item, suggesting that the model is unfamiliar with the concept. See A.4 for a deeper analysis of these cases. The remaining (1.3% of all lexical items) can be generated from the full lexical item, indicating that in these cases, the lexical item’s representation is distributed across multiple tokens. Focusing on the cases where at least one token represents the lexical item, we examine the number of representative tokens (those that result in an image that represents the item) and non-representative tokens (those that do not). We compare the proportion of representative tokens across different lexical items of different lengths. Our results (7 in A.3) show that typically one or two tokens represent the concept, while the remaining tokens are non-representative tokens. As the lexical item becomes longer, the number of non-representative tokens tokens increases. Focusing on these non-representative tokens, we find that they account for 52% of the tokens within the lexical items represented by least two tokens. In the next section, we examine the effect of removing non-representative tokens.
Figure 3: Examples illustrating the effect of removing non-representative tokens. Top row: Images generated after removing non-representative tokens (Representative tokens are shown in bold; non-representative tokens are in gray.). Bottom row: Images generated from the full prompt. Left: In most cases, removal results in no noticeable effect on the generation. Right: In some cases, removal improves alignment with the prompt.
4.2 Are all prompt tokens necessary?
Our analysis reveals that 52ֿof lexical item tokens are non-representative. In this section, we examine whether these tokens have any effect on image generation. To answer this question we apply our intervention method for each prompt, generating an image after masking all the non-representative tokens. We then use Qwen-VL to measure whether the image aligns with the prompt and compare it to the image generated without this intervention.
Non-representative tokens are redundant.
Our results show that removing non-representative tokens generally does not harm the generation (see 3, left-hand side). Of the cases where the original generation was aligned with the prompt, in 98% of the cases, the generated image still aligns with the prompt, suggesting that these tokens are largely redundant. Moreover, in cases where the original image failed to align with the prompt, we observe a 21% reduction in error rate (see 3, right-hand side). We attribute this improvement to the model being guided more by the representative tokens after removing the non-representative ones. See 2 in A.3 for full results.
Can we identify redundant tokens without generating their images?
Identifying which tokens are representative can improve image generation by removing redundant ones as shown in 3 (right). Moreover, many existing T2I methods operate at the token level, typically using all of a lexical item’s tokens to represent the item. (Chefer et al., 2023a; Dahary et al., 2024). Refining these methods to operate only on representative tokens could lead to improvements without requiring any additional changes. However, generating an image for each token is computationally expensive. To this end, we propose two ways to estimate token importance without image generation. First, we calculate the edit distance between a lexical item (e.g., “giraffe”, which is tokenized as “gir”, “a” and “ffe”, and a given token (e.g., ‘a’). We find that higher edit distance indicates a lower likelihood that the token is representative (Pearson correlation of $-0.44$). While this correlation is relatively strong, it is not accurate enough for practical use. Instead, we train a $k$-NN classifier ($k{=}5$, Euclidean distance) to predict token redundancy from encoded representations. We extract 6,966 tokens (4,864 unique lexical items) from our dataset, each labeled for whether the VLM labeled this token as representative. We find that this simple classifier achieves a precision of 92% in predicting token redundancy, highlighting its potential as an efficient tool for removing non-representative tokens while preserving the essential ones. Further details about the classifier can be found in A.1.
Figure 4: Evaluating inter-item information flow: Our proposed framework to interpret the information flow between lexical items in the prompt. For each lexical item, we generate an image from its contextual representations (left), and from it’s uncontextualized representation (right), and analyze the generated images using a VLM (middle). In this example, we interpret the item “pool” and assess whether it is influenced by the item “table”. To do so, we ask a VLM whether the image generated from the token “pool” contains a “table”. To ensure this is the result of information flow, and not a natural correlation between a pool and tables, we generate the item “pool” without context (right-hand side), and verify that “table” is not present in this image. If this is the case, we conclude that there was information flow from “table” to “pool”.
5 Inter-item information flow
User prompts are typically composed of multiple lexical items. Here we investigate the influence of lexical items on one another. For example, in the sentence “a pool next to a table”, is the representation of the item “pool” affected by the presence of the item “table”? More generally, given a prompt, made up of several lexical items, we aim to measure whether the presence in the prompt of one item affects the representation of another. To this end, we conduct the following experiment: given a prompt, we isolate each lexical item one at a time. First, we encode the entire prompt. Next, for each item, we patch the representations of all other tokens using our method Section 3.1, and generate an image based on the modified sequence. Specifically, we define a subset $S\subset\{1,\ldots,N\}$ containing the token indices of the selected lexical item, construct the patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$, and use it to guide the diffusion model. This process produces one image per lexical item in the prompt—each encoded in the full context of the prompt but with all other token representations masked. These images allow us to assess what information is encoded in each item’s representation. We then generate an image for the uncontextualized version of each lexical item—encoding only that item in isolation, without the rest of the prompt. To assess whether a lexical item $y$ influences another item $x$, we use Qwen2-VL to check if $y$ is present in the image generated from $x$, both with and without context. We then compare the Qwen2-VL answers between the contextualized and uncontextualized versions. Refer to A.1 for more details regarding Qwen2-VL and the prompts we use. An influence is considered True if $x$ did not exhibit $y$ in the uncontextualized representation but did in the contextualized one. Confirming that $y$ does not appear in the isolated representation of $x$ is crucial, as we want to ensure that the presence of $y$ in $x$ is due to information flow that occurs during textual encoding, and not due to some underlying correlation between the items. This approach allows us to quantify the degree to which lexical items affect one another within a prompt. See 4 for a demonstration of this framework.
Figure 5: Examples of information flow between items. Top: Images generated from a lexical item encoded alongside another item that alters its representation. Bottom: Images generated from the uncontextualized representation of the same lexical item. The first three images (from the left) demonstrate correct information flow, while the last image (far right) demonstrates incorrect information flow.
Our results show that in 89% of the cases, lexical items are not influenced by other items in the prompt. In the remaining 11%, one lexical item incorporates information about another item that appears in the same prompt; see 5 for examples. Interestingly, we find that when such influence occurs, it often involves items that frequently co-occur in natural language, and are therefore likely to appear together in the training data. For example, the item “pool” often shifts its representation from a swimming pool to a pool table when “table” is present in the context. This phenomenon occurs even when “table” is not used as a pool table, as in the sentence “A pool next to a table” (see 4). To measure the extent of information flow between items that should not influence each other (as in the “pool” and “table” example), we divide the identified cases of flow between items into two categories. The first includes cases where the items are semantically linked, such as “black” and “bear” in the prompt “a black bear”. The second includes cases where the items are not semantically linked, such as “black” and “bear” in the sentence “a green bear by the black tree”. To achieve this, we use an LLM as a judge (details in A.1). Our analysis reveals that 31% out of the influenced instances exhibit unintentional information flow.
Application.
Based on our findings, we present a method for mitigating semantic leakage. Our approach first encodes the prompt as usual, then separately encodes the item suspected of leakage. Finally, we replace the representation of the contextualized, leaked item in the original prompt’s encoding with its uncontextualized representation, and condition the diffusion process on this mixed representation (see 6). To evaluate the effectiveness of our method, we adapt prompts from Rassin et al. (2022) as these prompts often induce semantic leakage. For example, given the prompt “a bat is flying over a baseball stadium”, models tend to generate a flying baseball bat rather than the animal, due to the strong association between “bat” and “baseball stadium”. Since the dataset contains only about 30 examples, we use GPT-4o to expand it to 110 prompts. See A.2 for details on the augmentation process, including the full list of prompts. Out of the 110 prompts, 104 exhibit semantic leakage (as determined by manual human evaluation), where the image contains a concept not intended by the prompt. We then select the lexical items in each prompt that we suspect of having leaked, and apply our method on them. Remarkably, we achieve an 85% improvement on this set. See 8 in A.3 for additional qualitative examples.
Figure 6: Removing semantic leakage by replacing the contextually leaked concept representation. (1) Regular generation produces an image showing a crosswalk to the right of a bus station. (2) Generation from the prompt “standing zebra”, without any context, results in the correct interpretation of the zebra as an animal. (3) Generation using the original prompt, but with the leaked concept “standing zebra” replaced by it’s uncontextualized representation, yields a correct image.
6 Conclusions
In this study, we examined the information flow within lexical items and between different lexical items during the textual encoding of a prompt in T2I models. We introduced an intervention to evaluate how information is distributed across various token representations. For in-item information flow, our analysis revealed that information within a concept is not uniformly distributed; instead, it is primarily concentrated in a few specific tokens. For inter-item information flow, we found that information typically does not flow between different items. When it does, the flow is often unintended and may result in semantic leakage. Our experiments suggest that semantic leakage originates within the text encoder itself, challenging the common assumption that each token primarily captures its intended meaning. We showed that semantic leakage phenomenon can be largely mitigated by simply encoding the leaked item separately and patching it into the prompt representation before the diffusion. While prior work has focused on modifying cross-attention to address challenges in text-to-image models, our results suggest that these methods should also account for information flow between tokens that occurs earlier, during the textual encoding.
7 Acknowledgments
This research was supported in part by the Israel Science Foundation (grant No. 448/20), NSF-BSF grant 2020793, an Azrieli Foundation Early Career Faculty Fellowship, an AI Alignment grant from Open Philanthropy, the European Union (ERC, Control-LM, 101165402), and by an Academic Gift by NVIDIA. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Appendix
A.1 Technical details
Lexical item classification
We define a lexical item as either a single word or a compound expression of multiple words that, in context, conveys a unified semantic meaning. A compound expression is treated as a single item when its words form a fixed lexical unit with cohesive semantics rather than merely exhibiting a modifier–head relationship. For example, while “broken mirror” describes a mirror’s state, expressions like “hot air balloon” or “teddy bear” denote entities with distinct identities. Similarly, although phrases such as “identical twins” or “baseball bat” might be interpreted as separate concepts, conventional usage supports their treatment as unified entities. We employ the reasoning model 03-mini-high as a classifier to tag multi-word lexical items in both the target prompts and the dataset. The model returns a list of identified multi-word expressions, while the remaining untagged words are treated as individual lexical items.
Redundant token classification
We propose a probing classifier to predict whether a token is redundant (i.e., non-representative of its lexical item) using solely its encoded representation, without the need to generate an image. For this purpose, we extracted the 6,966 tokens corresponding to 4,864 unique lexical items from our dataset. Each token was annotated with a binary label indicating whether it represents the lexical item by the VLM (see A.1 for more details). We split the data into training and validation sets using an 80-20 ratio. A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations. The results on the evaluation set are presented in 1. The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
Table 1: Performance of k-NN classifier ($k=5$) for predicting token redundancy
Metric Accuracy Precision Recall F1-score
Score 0.82 0.92 0.74 0.82
Evaluation visual generations.
We evaluate whether an image matches a textual description using Qwen2-VL-72B-Instruct (Wang et al., 2024). The following prompt is used:
”In Yes, No and maybe. Does every image match one of those descriptions: (description string)? Answer Yes if all images match or relate to at least one description, Maybe if only some match, otherwise No.”
Here, the textual description can be either a single lexical item or a complete textual prompt.
Evaluating generated textual descriptions.
We employ gpt4o (OpenAI et al., 2024) to evaluate the textual interpretations produced by Patchscopes. We use the following prompt:
“In Yes, No and Maybe. Does every image match the description: {Patchscopes_description} ? Answer Yes if all images match or relate to the description, Maybe if only some match, otherwise No.”
Classifying the relations between items.
We enhance our leakage validation by distinguishing between cases where two lexical items exhibiting semantic leakage are perceptually bound together—for example, “old” and “man” in the prompt “a portrait of an old man”—and cases where they are not as “cone hat” and “eating” in the prompt “A person searing a cone hat is eating” (see 5). To achieve this, we use a large language model (LLM) as a judge. Specifically, we employ the following prompt:
“In Yes or No: in this prompt: {input_prompt}, are {item_1} and {item_2} perceptually bound together?”
We then filter out all cases where the lexical items are perceptually bound together and find that only 6.5% instances exhibit unintentional leakage.
A.2 Data
DrawBench (Saharia et al., 2022): We include all categories except for “misspelling”, “rare words”, and “text”. Overall we end up with 134 prompts from DrawBench. Parti Prompts (Yu et al., 2022a): We include all categories except “Style & Format”, “Writing & Symbols”, and “Arts”. Overall we end up with 923 prompts from Parti Prompts. In total, we end up with 1,056 prompts.
Extended dataset of leakage prompts
Our augmentation process incorporates two components. First, we generate variations of existing prompts from Rassin et al. (2022) (e.g., modifying ’a gentleman with a bow in the forest’ to ’a man wearing a bow in the jungle’). Second, we introduce novel prompts with potential semantic leakage. For these prompts, we applied a one-lexical item change test by generating an image from a similar prompt that substitutes the affected or leaked item with an alternative term (e.g., replacing ’bishops’ with ’cardinals’ or ’checkers’ in ’chess in ’2 bishops playing chess’). This test ensures that minimal lexical modifications do not alter the intended semantic meaning while producing a different image due to semantic leakage from another item in the prompt (see the first two columns in 8 for few visual examples) . Together, these methods enrich the dataset and provide a robust framework for analyzing semantic leakage. The full list of prompts is available in the following anonymous Git repository: https://anonymous.4open.science/r/TokenRole-3E19.
A.3 Additional results
Figure 7: Distribution of representative vs. non-representative tokens across different lexical item lengths.
Method Success Rate
Regular Generation 84.1%
After Removing Redundant Tokens 87.6%
Table 2: Image-text alignment success rate (evaluated with Qwen-VL): comparison between standard generation and generation with redundant tokens removed
Figure 8: Examples from our semantic leakage method. Left: standard generation of leakage contained prompt. Second: generation using a one-lexical item change test as part of the dataset creation (a minimal substitution to verify that a slight lexical change yields a different image). Third: image from the contextual representation (misinterpreted item). Forth: image from the uncontextualized representation (correct interpretation). Right: final generation after patching the correct, uncontextualized representation into the prompt.
A.4 Catastrophic negligence
Catastrophic negligence refers to cases where a concept or entity is completely absent from the generated image. Chefer et al. (2023a) attempted to address this issue by increasing attention to the neglected item, implicitly assuming that the model understands the concept. We find that in 7% of the cases, the item is not generated from any of its constituent tokens and is absent from the full prompt’s image. To better understand this phenomenon, we leverage Patchscopes (Ghandeharioun et al., 2024), a tool that interprets intermediate token representations by generating textual explanations using the decoder in encoder-decoder architectures. Using Patchscopes, we observe that in approximately 67% of these cases the tokens are explained coherently, indicating that the item is adequately represented in the text encoder. For example, when we patch the encoded representation of the lexical item tuba into the prompt “describe X”, Patchscopes describes it as “a musical instrument played by blowing into the mouthpiece.” However, when we try to generate an image from its representation, the model fails (see 9), suggesting that the diffusion model may have encountered too few examples of the intended entity during training. Additionally, we identify two abstract categories that consistently result in failure in image generation, while Patchscopes accurately describes the concept: negation and numbers. For instance, the lexical item “without” failed to generate the an image without the item in 100% of 100 observed instances. See  9.
Figure 9: Cases where intervention-based interpretations fail, while Patchscopes accurately describe token meaning. Each column corresponds to a different prompt. The first row shows images generated from the contextual bolded token. The second row presents the same token explained by Patchscopes, and the third row shows the final image generated from the full prompt. Left: The image generated from the token “tub” (from “tuba”) depicts a lamp, whereas Patchscopes correctly identifies it as a musical instrument. Middle: The concept “without” fails to yield an image lacking a refrigerator, although Patchscopes correctly describes its semantics. Right: The concept “7” leads to an incorrect number of entities in the image, while Patchscopes correctly identifies the quantity.
A.5 Additional models
A.5.1 FLUX-Dev
In addition to our primary experiments with FLUX, we repeated all analyses using the Flux-dev variant. The redundant versus representative token experiments yielded similar trends, with 55% of tokens identified as representative and 45% as non-representative—values closely matching those observed with FLUX. Likewise, our inter-item flow experiments confirmed that information flow occurred in 11% of cases (and 3.1% miss intended leakage), reinforcing the overall patterns reported in the main text. Notably, while the aggregate trends are consistent across models, the specific lexical items resolved can differ between FLUX-schnell and Flux-dev, indicating potential a slightly different inner-lexicon. These findings underscore the robustness of our approach while highlighting model-dependent nuances in token representation and information flow dynamics.
A.5.2 SDXL-Turbo
Figure 10: Images generated from individual subtokens in SDXL-Turbo. We find that, in many cases, the representation of an item is not clearly reflected in any of its subtokens—for example, in the case of the token “skateboard.” Another interesting observation is that the last token of a lexical item often carries its representation, as seen in the “square” token of “times square.” We also observe that the EOT token incorporates information from the full prompt.
Our analysis reveals that SDXL-Turbo, which uses the CLIP text encoder, behaves markedly differently from FLUX, which relies on the encoder in the encoder-decoder T5-XXL. In SDXL-Turbo, the text encoder is a causal language model, meaning each token’s encoding is influenced only by its preceding tokens during the encoding process. We repeated our in-item information flow experiments using SDXL-Turbo. Our first observation is that most generated images are either abstract or unrelated to the intended lexical items. According to our analysis, 55% of lexical items in SDXL-Turbo lack any representative token (compared to just 11% in FLUX). Moreover, when a representative token is present in CLIP, it is typically the final token of the lexical item (see 10). This is aligned with the unidirectional encoding of the model. Another phenomenon we observe—consistent with CLIP’s training objective—is the unusually dominant role of the end-of-sequence (EOS) token. Images generated from the EOS token often encapsulate nearly the full semantic content of the prompt. In our evaluation, 62% of EOS-generated images matched the prompt ( compared to 73% when using the full prompt). We believe this also causes our intervention method to be less effective, since when we interpret a single token, we patch all other tokens, including the EOT token—which usually contains a lot of information—with tokens derived from an empty prompt (see 10).

### v2

@@
 Abstract
-Text-to-Image (T2I) models often suffer from issues such as semantic leakage, incorrect feature binding, and omissions of key concepts in the generated image.
-This work studies these phenomena by looking into the role of information flow between textual token representations.
-To this end, we generate images by applying the diffusion component on a subset of contextual token representations in a given prompt and observe several interesting phenomena.
-First, in many cases, a word or multiword expression is fully represented by one or two tokens, while other tokens are redundant.
-For example, in “San Francisco’s Golden Gate Bridge”, the token “gate” alone captures the full expression.
-We demonstrate the redundancy of these tokens by removing them after textual encoding and generating an image from the resulting representation.
-Surprisingly, we find that this process not only maintains image generation performance, but also reduces errors by 21% compared to standard generation.
-We then show that information can also flow between different expressions in a sentence, which often leads to semantic leakage.
-Based on this observation, we propose a simple, training-free method to mitigate semantic leakage: replacing the leaked item’s representation after the textual encoding with its uncontextualized representation.
-Remarkably, this simple approach reduces semantic leakage by 85%.
-Overall, our work provides a comprehensive analysis of information flow across textual tokens in T2I models, offering both novel insights and practical benefits.
-[^1] We release our code at https://github.com/tokeron/lens
+Text-to-image (T2I) models generate images by encoding text prompts into token representations, which then guide the diffusion process.
+While prior work has largely focused on improving alignment by refining the diffusion process, we focus on the textual encoding stage.
+Specifically, we investigate how semantic information is distributed across token representations within and between lexical items (i.e., words or expressions conveying a single concept) in the prompt.
+We analyze information flow at two levels: (1) in-item representation—whether individual tokens represent their lexical item, and (2) cross-item interaction—whether information flows across the tokens of different lexical items.
+We use patching techniques to uncover surprising encoding patterns.
+We find information is usually concentrated in only one or two of the item’s tokens—For example, in the item “San Francisco’s Golden Gate Bridge”, the token “Gate” sufficiently captures the entire expression while the other tokens could effectively be discarded.
+Lexical items also tend to remain isolated; for instance, the token “dog” encodes no visual information about “green” in the prompt “a green dog”.
+However, in some cases, items do influence each other’s representation, often leading to misinterpretations—e.g., in the prompt “a pool by a table”, the token pool represents a pool table after contextualization.
+Our findings highlight the critical role of token-level encoding in image generation, suggesting that misalignment issues may originate already during the textual encoding.
+Follow the Flow: On Information Flow Across Textual Tokens in Text-to-Image Models
+[…]
+Figure 1: Our main findings.
+Left: Information within a lexical item is unevenly distributed across its tokens’ contextualized representations.
+In this example, one token carries the meaning of the entire item (e.g., lic represents a pelican, while pe and an do not).
+Right: Context may cause incorrect information flow between items.
+In the example, the representation of runway is influenced by the context of “a businesswoman” (vs.
+“businessman”).
 1 Introduction
-Text-to-image (T2I) models consist of two main components: a text encoder and a diffusion model (Ho et al., 2020; Song & Ermon, 2019).
+Text-to-image (T2I) models typically consist of two main components: a text encoder and a diffusion model (Ho et al., 2020; Song and Ermon, 2019).
 The former processes the user’s prompt, transforming it into a representation that guides the latter in generating the image.
-While T2I models are powerful and widely used, misalignment issues frequently arise.
-Generated images often fail to capture key concepts from the user’s prompt, leading to catastrophic neglect (Chefer et al., 2023a), semantic leakage of properties between entities (Rassin et al., 2022), or incorrectly bind different concepts (Huang et al., 2023; Rassin et al., 2022).
-Prior work has attempted to address these issues by modifying the cross-attention mechanism between the diffusion model and the encoded textual tokens (Rassin et al., 2023; Chefer et al., 2023a; Dahary et al., 2024), under the implicit assumption that the text encoder accurately captures the user’s intent and that each textual token primarily represents the item it is intended to convey.
-In this work, we challenge these assumptions and investigate how information flows within tokens during the textual encoding process, and what information is encoded in each textual token representation.
+Though widely used, T2I models often exhibit prompt-image misalignment, where generated images fail to capture key concepts from the user’s prompt (Chefer et al., 2023a; Rassin et al., 2022; Huang et al., 2023a).
+Prior work has attempted to address these issues by modifying the diffusion stage, and particularly the cross-attention mechanism (Rassin et al., 2023; Chefer et al., 2023a; Dahary et al., 2024), under the implicit assumption that each textual token reliably encodes the item it is intended to convey.
+This raises two fundamental questions: First, is the information for an intended concept distributed evenly across the concept tokens after textual encoding, or concentrated in just a few tokens?
+Second, does each token exclusively encode that concept?
+In this work, we examine this assumption and study how visual information is distributed across tokens during the textual encoding stage.
+We focus our analysis on lexical items—words or phrases that convey a single concept.
+We trace how item information is distributed both within the tokens of a single item (in-item), and across tokens of different items (cross-item) (see Fig.˜1 for examples of the different cases).
+Using a causal intervention framework (Toker et al., 2025), we systematically assess what information is encoded in each contextual token representation at the encoder’s output, using prompts from widely used T2I benchmarks as our evaluation setting (Section˜2.2).
 [^1] Throughout this work, we study token representations at the output of the text encoder.
 For brevity, we sometimes omit the word “representation”.
-To this end, we analyze lexical items within prompts, where each lexical item is a single word or multiword expression that functions as a unit of meaning.
-Our analysis focuses on two complementary aspects.
-First, in-item information flow—how different sub-word token representations of the same lexical item encode and distribute information.
-For example, do all the token representations of the item “pelican” (“pe”, “lic”, and “an”) encode this lexical item (Fig.
-1, top left).
-Second, inter-item information flow—how information flows between different lexical items (Feng et al., 2024; Feng & Steinhardt, 2024).
-For example, given the prompt “A businesswoman on a runway”, does the representation of “runway” change in the context of the word “businesswoman” (Fig.
-1, bottom middle).
-Figure 1: Our main findings.
-Left: Information within a lexical item is unevenly distributed across its token representations, with a few carrying most of the meaning.
-Middle: Context may cause incorrect information flow between items, distorting their representations and leading to unintended images.
-In this example, the representation of “runway” is distorted in the context of “a businesswoman”.
-Right: We propose a simple solution—encode the leaked item (“runway” in this case) separately and patch it into the representation of the prompt after textual encoding.
-For in-item information flow (4), we adapt a causal intervention method (Toker et al., 2025; see Section 3.2) to generate images from individual token representations.
-Specifically, we first encode the entire sentence, and then generate an image from each individual contextualized token representation.
-The resulting image reflects the information that the diffusion model can reconstruct when conditioned on that single token’s contextual representation.
-Our experiments with FLUX (black-forest labs, 2024) show that, in most cases, each lexical item is fully represented by only one or two tokens, while its remaining tokens might be considered redundant (4.2).
-Finally, we demonstrate that our findings can be used to improve image–text alignment in image generation.
-To this end, we mask all the non-representative tokens (tokens that do not represent the lexical item) just before the diffusion, and generate images directly from the remaining representations.
-We find that this operation alone reduces image generation error by 21% compared to standard image generation.
-For inter-item information flow (5), we adapt the same causal intervention method—this time to assess the information encoded in one lexical item with respect to others in the prompt.
-For instance, in the sentence a “A pool and a table”, we examine the information flow between the items “pool” and “table”.
-We find that in 89% of cases, each lexical item’s information is preserved and not influenced by others.
-However, in some cases, information flows between tokens.
-Interestingly, this flow can occur between lexical items that should not influence each other, for example, “table” influencing the representation of “pool” in the sentence above, changing pool’s representation to a “pool table”.
-We show that this leakage can be mitigated by patching the leaked lexical item’s representation with an uncontextualized representation of this item.
-Notably, this simple intervention prevents leakage in 85% of cases.
-In summary, we analyze information flow in T2I models and uncover two key findings: (1) lexical items are often represented by just one or two tokens, and removing non-representative tokens can improve generation; (2) while different lexical items usually do not influence each other’s representations, when they do, it can happen between items that should not influence each other—leading to semantic leakage.
-Finally, based on our observation, we propose a simple method to mitigate semantic leakage, leading to a 85% reduction in leakage errors.
-2 Related work
-2.1 Interpretability
-Interpretability in multimodal models.
-Previous work on interpretability in T2I models analyzed the text encoder.
-Chefer et al.
-(2023b) examined the concepts encoded in the latent space of CLIP (Radford et al., 2021) and how different concepts relate to each other.
-Toker et al.
-(2024) studied the progression of textual representations through the layers of the text encoder, using the diffusion model as a lens into the encoding process.
-Another line of work explored the connection between the text encoder and the image generator (Tang et al., 2023).
-In this work we investigate how information flows between tokens within a single layer.
-We focus on the final layer, as it directly conditions the diffusion process.
-Information flow in LLMs.
-Interpretability research has examined how token information propagates via attention in LLMs (Vig & Belinkov, 2019; Clark et al., 2019).
-However, raw attention weights can be misleading (Pruthi et al., 2020; Jain & Wallace, 2019).
-To address this, methods such as Attention Rollout (Abnar & Zuidema, 2020) were developed to more reliably measure information flow through attention.
-Other studies have employed Patchscopes (Ghandeharioun et al., 2024) and the logit lens (nostalgebraist, 2020) to interpret the encoding processes in LLMs.
-These works demonstrate that modern language models go far beyond simple subword processing.
-For instance, Kaplan et al.
-(2025) showed that models perform an implicit detokenization process, effectively fusing subword tokens into coherent word-level representations.
-Feucht et al.
-(2024) reported an “erasure” effect, in which the contributions of earlier tokens are overwritten by later ones—suggesting the emergence of an internal lexicon.
-Another prominent direction is probing (Adi et al., 2016; Liu et al., 2019).
-For example, Zhang & Bowman (2018); Brunner et al.
-(2019) trained probes to classify, from a token’s intermediate representation, the identity of tokens a fixed number of positions away.
-However, these methods are not fully reliable, as probes can exploit spurious correlations (Belinkov, 2022).
-Moreover, these works typically assess whether information is present or flows between tokens, but not whether the next step in the pipeline (e.g., the LM head or decoder) actually uses this information.
-In our analysis, we focus on intervention methods that test whether the next component in the image generation pipeline can retrieve and use the relevant information.
-2.2 Challenges and solutions in multimodal models.
-The recent rise of T2I models has introduced several challenges, prompting various approaches to address them.
-This work shows that a better understanding of information flow between tokens can help mitigate many of these challenges.
-Semantic leakage.
-One notable challenge in T2I models is semantic leakage (Rassin et al., 2022).
-For example, given the prompt “a bat flying in a baseball stadium”, the generated image might include a baseball bat instead of a bat (the animal).
-A similar effect has been observed in language models, where the context influences generation even when it conceptually should not (Gonen et al., 2024).
-Attribute binding.
-Another issue is attribute binding, which concerns the correct association of attributes with entities mentioned in the prompt.
-For example, in the prompt “a yellow flamingo and a pink sunflower”, models must correctly bind “yellow” to “flamingo” and “pink” to “sunflower”.
-Several approaches have been proposed to address these challenges, including performing an optimization in the latent space to produce an attention pattern that is better aligned with the syntactic structure of the sentence (Rassin et al., 2023), and merging tokens to concentrate information into a single representation (Hu et al., 2024).
-Catastrophic neglect.
-A further prominent challenge is catastrophic neglect, where entities mentioned in the prompt are completely omitted from the generated image.
-Many works have attempted to mitigate this, one of the first being Attend and Excite (Chefer et al., 2023a), which optimizes the image latents to increase attention on the neglected entity tokens.
+For in-item representation (Section˜3), we find that a lexical item’s meaning is typically concentrated in one or two representative tokens.
+Surprisingly, ablating the non-representative tokens reduces generation error by 21% in FLUX (black-forest labs, 2024).
+In contrast, when no representative tokens exist for a particular item, it is often omitted entirely from the image—an instance of item negligence (Chang et al., 2024a; Chefer et al., 2023a).
+For cross-item interactions (Section˜4), we assess whether information flows between different lexical items in a prompt.
+We observe such cross-item flow in 11% of cases.
+Interestingly, the flow of information does not always follow sentence syntax, which can result in misrepresentation of items within specific contexts—especially with polysemous words.
+For instance, in the prompt “a pool by a table”, “pool” can wrongly suggest “table” refers to a billiard table.
+In summary, we study the textual encoding and identify two main insights.
+First, Lexical items have concentrated representation—the meaning of the item is often concentrated in a single representative token.
+Second, while items in a sentence do not typically encode information regarding other items in the sentence, in some cases, often when polysemous are involved, semantic leakage (Dahary et al., 2024; Gonen et al., 2024; Rassin et al., 2022) occurs during the textual encoding phase, causing miss-interpretations of items.
+Together, these insights suggest practical avenues for improving token-level intervention techniques, enhancing evaluation benchmarks, and guiding future encoder-aware T2I generation approaches.
 2 Methodology
 2.1 Intervention on the Text Encoder
-Our goal in this paper is to evaluate the effect of information flow between textual token representations, which later condition the diffusion process.
-We are interested in analyzing how information is distributed across tokens within a lexical item (e.g., do all the tokens in “San Francisco’s Golden Gate Bridge” represent that item, or is the information more diffusely spread?).
-To do so, we must be able to interpret the information encoded in individual tokens.
-We are also interested in measuring how different lexical items influence one another.
-To this end, we seek to isolate a subset of tokens and evaluate their joint representation.
-For these purposes, we adjust text intervention in diffusion models (Toker et al., 2025), allowing us to generate images from arbitrary subsets of tokens within the encoded user prompt.
-Given a prompt with $N$ tokens, $t_{1},t_{2},\ldots,t_{N}$, our goal is to isolate and interpret the information encoded by a subset of these tokens.
-Let $S\subset\{1,\ldots,N\}$ be the index set of selected tokens, where $1\leq|S|\leq N$.
-We begin by encoding the full prompt using the text encoder $E$, yielding the final hidden states $e_{1},\ldots,e_{N}$.
+Our goal in this paper is to evaluate the effect of information flow between textual token representations, that subsequently condition the diffusion process.
+In Section˜3, we are interested in analyzing how information is distributed across tokens within a lexical item .
+To do so, we seek to interpret the information encoded in individual tokens.
+In Section˜4, we are interested in measuring how different lexical items influence one another.
+To this end, we seek to isolate a subset of tokens and evaluate their joint representation in the context of another item.
+For these purposes, we adjust a method for text intervention in diffusion models proposed by Toker et al.
+(2025), allowing us to generate images from arbitrary subsets of contextual token representations by masking the rest of the tokens in the sequence.
+Figure 2: Evaluating in-item information flow.
+Our proposed framework interprets the information flow within a lexical item.
+We generate images from each token comprising the lexical item (left) and analyze them with a VLM (right).
+In this example, only the token lic represents the concept “pelican”, whereas pe and an do not.
+Given a prompt with $N$ tokens $t_{1},t_{2},\ldots,t_{N}$, our goal is to isolate and interpret the information encoded by a subset of these tokens.
+Let $S\subset\{1,\ldots,N\}$ be the index set of selected tokens, where $0<|S|<N$.
+We begin by encoding the full prompt using the text encoder $E$, yielding the final hidden states $h_{1},\ldots,h_{N}$.
 Separately, we encode a sequence consisting entirely of pad tokens to obtain pad embeddings $p_{1},\ldots,p_{N}$.
 We then construct a patched prompt by replacing all hidden states outside $S$ with the corresponding pad embeddings:
-$\tilde{t}_{i}=\begin{cases}e_{i}&\text{if }i\in S,\\ p_{i}&\text{otherwise}\end{cases}\quad\text{for }i=1,\ldots,N.$
+$\tilde{t}_{i}=\begin{cases}h_{i}&\text{if }i\in S,\\ p_{i}&\text{otherwise}\end{cases}\quad\text{for }i=1,\ldots,N.$
 The patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ is then used to guide the diffusion model.
 Generating an image from this patched representation allows us to visualize and isolate the individual contributions of the selected tokens as interpreted by the diffusion model.
-To evaluate the information in the generated images, we use a vision-language model (VLM): For the inter-item information flow experiments, we assess whether the generated image represents the full lexical item, while for the in-item information flow experiments, we measure whether information from other lexical items is present in the generated image.
-See Section 3.2 for an illustration of the method applied to interpret individual subtoken representations.
+To evaluate the information in the generated images, we use a vision-language model (VLM): For the in-item representation experiments, we assess whether the generated image represents the full lexical item, while for the cross-item interactions experiments, we measure whether information from other lexical items is present in the generated image.
+See Fig.˜2 for an illustration of the method on the  in-item representation  experiment.
 2.2 Experimental Setup
 Models.
-We experiment with FLUX-Schnell (black-forest labs, 2024)—a recent state-of-the-art T2I model.
-We also report similar findings on FLUX-dev in A.5.1.
+Our main results are reported for FLUX-schnell (black-forest labs, 2024), a recent state-of-the-art T2I model.
+We repeat all experiments with FLUX-dev, SDXL-Turbo (Sauer et al., 2023) and SANA (Xie et al., 2024a).
 FLUX models employ T5-XXL (Raffel et al., 2019) as their text encoder, enabling bidirectional information flow between tokens.
-In A.5.2 we also include another model, SDXL-Turbo (Sauer et al., 2023), which uses a different text encoder—CLIP (Radford et al., 2021).
-Unlike T5-xxl, CLIP is a causal language model.
-See Appendix A.5.2 for the analysis of SDXL-Turbo.
+Unlike Flux, SDXL used CLIP (Radford et al., 2021), and SANA uses Gemma (Team et al., 2024).
+These text encoders are unidirectional, meaning the tokens are only influence by the tokens that came before.
+These changes cause some differences in the information flow between tokens, elaborated in Section˜6 and in Section˜A.3.2
 Data.
-We use a subset of prompts from DrawBench (Saharia et al., 2022) and Parti (Yu et al., 2022b).
-In total, we use 1,053 prompts.
+We use a subset of 1,053 prompts from DrawBench (Saharia et al., 2022) and PartiPrompts (Yu et al., 2022b) datasets, filtered to include 4–20 words prompts and exclude cases with added complexity unrelated to our focus (e.g., misspellings, written text, rare words).
 For each prompt, we generate five images using different random seeds.
-To conduct the analysis in our paper, we first find the lexical items in each prompt.
-We use GPT-4o (OpenAI et al., 2024) to extract the lexical items, resulting in a total of 4,864 unique lexical items.
-See A.1 for the exact prompt we used and further technical details.
+To extract the lexical items, we prompt GPT-4o (OpenAI et al., 2024), resulting in 4,864 unique items across prompts.
+See Section˜A.1.1 for further technical details.
 We then use spaCy (Honnibal et al., 2020) to determine the part of speech of each lexical item, and retain only nouns, proper nouns, and adjectives, as these are typically concrete and can be identified in their visual representation.
-We end up with 3,891 unique lexical items and use them to evaluate both inter-item information flow and in-item information flow.
+We end up with 3,891 unique lexical items and use them to evaluate both in-item representation and cross-item interactions.
 Evaluation.
-To evaluate the content of generated images, we also employ Qwen2-VL-72B-Instruct (Wang et al., 2024), a model with strong general vision capabilities.
-We restrict our evaluation to binary (yes/no) questions.
-See A.1 for the exact prompt used and additional details.
-Figure 2: –ֿ
-Evaluating in-item information flow: Our proposed framework interprets the information flow within a lexical item.
-We generate images from each token comprising the lexical item (left-hand side) and analyze the resulting images using a VLM (right-hand side).
-In this example, we interpret the different tokens composing the item pelican and find that only one of them, lic, represents the concept of a pelican, while the other two tokens (“pe”, and “an”) do not.
-4 In-item information flow
+To evaluate the content of generated images, we employ Qwen2-VL-72B-Instruct (Wang et al., 2024), a model with strong general vision capabilities.
+We restrict our evaluation to binary (yes/no) questions and use the model to assess prompt-image alignment and the presence of specific items in the image (see Section˜A.1.3 for additional details).
+To validate Qwen2-VL’s reliability, we also conducted a human evaluation on 100 randomly sampled cases, evenly split between in-item representation (Section˜3) and cross-item interactions (Section˜4) settings.
+The results indicate substantial agreement, with a Cohen’s Kappa values of 0.868 for in-item representation and 0.764 for cross-item interactions.
+Next, we determined the majority agreement among the human annotators and compared it to the model predictions.
+For cross-item interactions, the accuracy (F1) is 0.927 (0.933), and for cross-item interactions, the accuracy is 0.810 (0.740).
+These results suggest that Qwen2-VL predictions align well with human judgment for our tasks.
+Acc. Relative %
+Tokens
+Removed
+# Prompts Before After Unaffected Degraded Improved Accuracy $\Delta$ ($p$)
+1 144 81.25 83.33 98.29 0.85 14.83 +2.08
+2 98 82.65 88.78 100.00 0.00 35.27 +6.13
+3 45 93.33 93.33 97.62 2.38 33.28 +0.00
+4 24 87.50 91.67 100.00 0.00 33.36 +4.17
+5+ 28 78.57 85.71 100.00 0.00 33.32 +7.14
+Overall 339 83.48 87.02 98.90 0.71 25.00 +3.54
+Table 1: Effect of removing non-representative tokens on prompt-level accuracy.
+The table reports: (i) accuracy before and after token removal, (ii) the percentage of originally successful prompts that remained successful (Unaffected) or became failures (Degraded), and (iii) the percentage of originally failed prompts that were corrected (Improved).
+Overall, removing non-representative tokens rarely harms generation and often improves it.
+3 In-Item Representation
+In this section, we analyze information flow within lexical items at the token level.
+We first show semantic information is unevenly distributed across tokens, with typically one or two representing the item (3.1).
+We then show non-representative tokens are largely redundant and may even harm generation (3.2).
+Finally, we examine items neglected from the generated image, tracing failures to poor textual encoding or gaps in visual grounding (3.3).
 3.1 How is Information Distributed Across Tokens?
-In this section, we explore how information is distributed across the token representations within a lexical item.
-We aim to answer two main questions: Do all tokens encode the same information?
-And is the information evenly distributed across tokens or concentrated in specific token representations?
-These questions are central, as many T2I applications (Chefer et al., 2023a; Rassin et al., 2023; Dahary et al., 2024) treat all tokens of a lexical item equally.
-Discovering a non-uniform distribution of information across tokens could improve application effectiveness by focusing on the most informative tokens.
-We apply our intervention method (Section 3.1) to the hidden representation of each token in the prompt.
-We then generate an image conditioned on each token’s contextualized representation.
-We use Qwen2-VL to assess whether an image represents the lexical item it is part of.
-We repeat this analysis for each lexical item, for each prompt in our combined dataset (Section 3.2).
-The results show that in 89% of the cases, at least one of the tokens fully represents the lexical item.
-Out of the remaining 11%, about 85% cannot be generated even from the full lexical item, suggesting that the model is unfamiliar with the concept.
-See A.4 for a deeper analysis of these cases.
-The remaining (1.3% of all lexical items) can be generated from the full lexical item, indicating that in these cases, the lexical item’s representation is distributed across multiple tokens.
-Focusing on the cases where at least one token represents the lexical item, we examine the number of representative tokens (those that result in an image that represents the item) and non-representative tokens (those that do not).
-We compare the proportion of representative tokens across different lexical items of different lengths.
-Our results (7 in A.3) show that typically one or two tokens represent the concept, while the remaining tokens are non-representative tokens.
-As the lexical item becomes longer, the number of non-representative tokens tokens increases.
-Focusing on these non-representative tokens, we find that they account for 52% of the tokens within the lexical items represented by least two tokens.
-In the next section, we examine the effect of removing non-representative tokens.
+We begin by exploring how information is distributed across the token representations within a lexical item.
+We focus on two key questions: Do all tokens encode the same semantic information?
+And is the information evenly distributed across tokens, or rather concentrated in specific tokens?
+These issues are crucial as many T2I applications (Chefer et al., 2023a; Rassin et al., 2023; Dahary et al., 2024) treat all tokens of a lexical item in a given prompt equally.
+Uncovering an asymmetrical distribution of information could improve the effectiveness of such applications by focusing on the most informative tokens.
+Given a prompt and a lexical item, we feed the prompt to the text encoder, obtaining contextualized token representations.
+We then identify the tokens comprising the item and apply our intervention method (Section˜2.1) to generate an image conditioned on each token’s contextualized representation (i.e., without letting other tokens in the prompt influence the diffusion process).
+Finally, we use Qwen2-VL to assess whether the image represents the overall lexical item it is part of.
+We define a representative token as one whose isolated representation results in an image containing its corresponding lexical item.
+Tokens that do not meet this criterion are considered non-representative.
+We repeat this analysis for each lexical item and for each prompt in our dataset (Section˜2.2).
+Our results (Fig.˜3, bottom row) show that in most cases (89%), there is (at least) one representative token.
+Interestingly, in the remaining instances where no representative tokens exist, we find the item is also absent from an image generated from the full item (allowing all of the item’s tokens to guide generation) in 88% of instances.
+Figure 3: Distribution of representative tokens per lexical item length.
+“Rep.” denotes the number of representative tokens; the bottom row aggregates over all lexical items.
+In most cases, one or two tokens sufficient to represent the entire lexical item.
+As item length increases, the number of non-representative tokens grows accordingly.
+We next focus on instances where at least one token represents the lexical item, and examine the number of representative- and non-representative tokens across lexical items of different lengths, averaged across items.
+Our results (Fig.˜3, top rows) show that typically one or two tokens represent the concept, while the remaining tokens are non-representative.
+Further, as the token length of the lexical item becomes longer, the number of non-representative tokens increases (Fig.˜3, rightmost column).
+Inspecting all lexical items in our data composed of two or more tokens, these non-representative tokens account for 52% of their tokens.
+Therefore, we next examine the effect of removing non-representative tokens altogether.
 Figure 4: Examples illustrating the effect of removing non-representative tokens.
@@
 Right: In some cases, removal improves alignment with the prompt.
-4.2 Are all prompt tokens necessary?
-Our analysis reveals that 52ֿof lexical item tokens are non-representative.
-In this section, we examine whether these tokens have any effect on image generation.
-To answer this question we apply our intervention method for each prompt, generating an image after masking all the non-representative tokens.
-We then use Qwen-VL to measure whether the image aligns with the prompt and compare it to the image generated without this intervention.
+3.2 Tokens: Are They All Necessary?
+We now examine whether non-representative tokens—tokens that do not result in an image representing the item—have any effect on image generation.
+To answer this question, we apply our intervention method to each prompt, this time generating an image after masking all the non-representative tokens.
+We use Qwen-VL to measure whether this image aligns with the prompt, and compare it to an image generated from the full prompt without intervention.
 Non-representative tokens are redundant.
-Our results show that removing non-representative tokens generally does not harm the generation (see 3, left-hand side).
-Of the cases where the original generation was aligned with the prompt, in 98% of the cases, the generated image still aligns with the prompt, suggesting that these tokens are largely redundant.
-Moreover, in cases where the original image failed to align with the prompt, we observe a 21% reduction in error rate (see 3, right-hand side).
-We attribute this improvement to the model being guided more by the representative tokens after removing the non-representative ones.
-See 2 in A.3 for full results.
-Can we identify redundant tokens without generating their images?
-Identifying which tokens are representative can improve image generation by removing redundant ones as shown in 3 (right).
-Moreover, many existing T2I methods operate at the token level, typically using all of a lexical item’s tokens to represent the item.
-(Chefer et al., 2023a; Dahary et al., 2024).
-Refining these methods to operate only on representative tokens could lead to improvements without requiring any additional changes.
-However, generating an image for each token is computationally expensive.
-To this end, we propose two ways to estimate token importance without image generation.
-First, we calculate the edit distance between a lexical item (e.g., “giraffe”, which is tokenized as “gir”, “a” and “ffe”, and a given token (e.g., ‘a’).
-We find that higher edit distance indicates a lower likelihood that the token is representative (Pearson correlation of $-0.44$).
-While this correlation is relatively strong, it is not accurate enough for practical use.
-Instead, we train a $k$-NN classifier ($k{=}5$, Euclidean distance) to predict token redundancy from encoded representations.
-We extract 6,966 tokens (4,864 unique lexical items) from our dataset, each labeled for whether the VLM labeled this token as representative.
-We find that this simple classifier achieves a precision of 92% in predicting token redundancy, highlighting its potential as an efficient tool for removing non-representative tokens while preserving the essential ones.
-Further details about the classifier can be found in A.1.
+Our results (Table˜1) show that removing non-representative tokens generally does not harm generation (see Fig.˜4, left-hand side).
+When the original generation is aligned with the prompt, the generated image after non-representative token removal remains aligned in 98% of cases, suggesting these tokens are largely redundant.
+Surprisingly, in cases where the original image fails to align with the prompt, we observe a 21% improvement in alignment after removing non-representative tokens (see Fig.˜4, right-hand side).
+We attribute this improvement to the model relying exclusively on the remaining representative tokens, which encode the correct semantics of the item.
+[^2] We measure improvement rates with Qwen2-VL before and after masking non-representative tokens.
+See Section A.1.3 for further details.
+Figure 5: Comparing Patchscopes to image generation with token-level patching in cases of item negligence.
+We assess whether a token’s concept is preserved by comparing images generated from its contextual representation, Patchscopes’ textual interpretation, and the full prompt image.
+Left: “tub” (from “tuba”) is correctly described by Patchscopes but fails to ground visually.
+Right: “paw” (from “pawn”) has missing details by both interpretations, suggesting a gap in the encoder’s conceptual knowledge.
+3.3 Item Negligence
+So far, we have explored scenarios where an item lacks a representative token, often leading to its omission from the final generated output, i.e., item negligence (Chang et al., 2024b; Chefer et al., 2023a).
+We next investigate why this omission occurs.
+We examine two potential causes: either the text encoder poorly represents the concept; or the concept is well-encoded, but the diffusion model isn’t familiar with its visual appearance.
+To study an item’s encoding without relying on the diffusion model, we use Patchscopes (Ghandeharioun et al., 2024), which decodes token representations into their natural language descriptions, using the full encoder-decoder architecture of T5 for text generation.
+Specifically, we patch an item’s encoding into the template “describe <item>” and evaluate whether the output indeed describes the item using GPT-4o (OpenAI et al.
+2024; see Section˜A.1.4 for prompt details).
+We apply this process to all neglected lexical items found in Section˜3.1, and examine whether the encoder is familiar with these items.
+We observe that in 67% of all instances of negligence, Patchscopes returns correct descriptions (for example, as demonstrated in Fig.˜5, “tuba” yields “a musical instrument played by blowing into the mouthpiece”).
+This suggests that, in such cases, item negligence is caused by a failure in the diffusion model.
+In the remaining 33%, the Patchscopes outputs are incorrect, indicating a gap in the encoder’s semantic understanding or knowledge.
+This points to two distinct sources of item-level negligence: Representational gaps within the text encoder itself—consistent with findings that augmenting text-side item descriptions improves factual generation (Huang et al., 2025)—and visual grounding failures despite sufficient text encodings.
+Category Count Percentage
+# pairs 15,950 100.00%
+No Information Flow 14,251 89.35%
+Information Flow 1,699 10.65%
+   – Source before Reference 0835 49.15%
+   – Reference before Source 0823 48.44%
+Table 2: Distribution of information flow between lexical item pairs. Subcategories under “With Information Flow” indicate the order of source and reference.
+4 Cross-Item Interactions
+In Section˜3, we demonstrated that a single token can effectively encapsulate the semantics of an entire lexical item.
+This raises a broader question: What defines the segmentation boundaries of a lexical item?
+Can its tokens further encode adjacent items, or even the general surrounding context?
+To this end, we isolate each lexical item in a prompt and assess whether it encodes information about other items in the prompt.
+For each item, we generate images from its contextualized representation (encoded within the full prompt), and its uncontextualized representation (using the item’s text by itself as a prompt).
+We then use Qwen2-VL to evaluate whether any other item from the prompt appears in the contextualized image, but not in the uncontextualized one—indicating information flow introduced by the surrounding context.
+See Section˜A.4 for implementation details.
+Our results (Table˜2) show that in 89% of the cases, lexical items do not encode information from other items in the prompt.
+In the remaining 11%, one lexical item incorporates information about another item (see Fig.˜6 for qualitative examples).Interestingly, information flow can emerge between items with no direct syntactic relation in the prompt.
+To better understand the nature of information flow, we analyzed the cases where lexical items appeared to influence one another.
+We categorize these prompt into 2 groups, based on whether this influence aligns with the prompt’s syntactic structure or not.
+The first category involves items with a direct syntactic relationship, such as the adjective “black” modifying the noun “bear” in the prompt “a black bear”.
+The second category includes items that are syntactically unrelated, such as the adjective “red” influencing the noun “building” in “a red car next to a building”.
+Using GPT-4o to classify between these cases (see Appendix A.1.5), we find that 31% of these information flow instances occur between syntactically unrelated items.
+Figure 6: Examples of information flow between items.
+Top: Images generated from a lexical item encoded alongside another item that alters its representation.
+Bottom: Images generated from the uncontextualized representation of the same lexical item.
+The first three images (from the left) demonstrate correct information flow, while the last image (far right) demonstrates incorrect information flow.
+Wrong information flow leads to misinterpretation.
+Our analysis reveals a recurring theme of semantic influence between items within a prompt, a pattern not prominent in most of the prompts.
+In these cases, the interpretation of one item is skewed by another semantically related, but syntactically distant, item.
+For instance, in the prompt “A standing zebra to the right of a city bus station”, the strong semantic connection between “bus station” and “zebra” leads the model to generate a “zebra crossing” instead of the animal (see Fig.˜7).
+We refer to this outcome, where contextual association overrides the prompt-intended meaning, as an incorrect item resolution.
+We note that such failures typically occur in cases of polysemous words.
+For example, the item bats is interpreted as wooden baseball bats in the prompt “bats fly around a baseball stadium”, whereas the intended meaning is obviously the flying animal.
+In such cases, the context overrides the correct sense of the item, leading to inaccurate generation.
+We hypothesize that these cases of incorrect item resolution are caused by information flow from the context to the misinterpreted lexical item.
+To validate this, We use the dataset from Rassin et al.
+(2022), which includes prompts known to induce semantic leakage due to implicit lexical associations (e.g., “bat” and “baseball stadium”).
+Unlike standard T2I benchmarks such as Huang et al.
+(2023b), which focus on spatial or visual challenges, these prompts highlight failures rooted in the encoding process itself.
+Since the original set includes only 30 examples, we expand it to 110 prompts using GPT-4o.
+See Section˜A.2 for augmentation details and the full prompt list.
+To evaluate our hypothesis, we conduct a simple causal test.
+We generate three images for each prompt: an image from the full prompt; another from the item without any context; and one from the full prompt, but with the lexical item’s tokens replaced by their uncontextualized representations (see Fig.˜7).
+For each image, we assess if the image depicts the intended interpretation of the item within the prompt (see A.1.6 for more details on this evaluation process).
+Our findings show that in 93% of the evaluated prompts, the item is misinterpreted in the original full-prompt image but correctly interpreted in the patched versions.
+This suggests that the failure stems from incorrect item resolution during the textual encoding phase: the language model misrepresents the intended meaning of the item due to contextual interference, retrieving its incorrect sense.
+Importantly, this indicates a failure in semantic resolution that originates entirely within the text encoder, and cannot be mitigated during image generation.
+Addressing such failures requires solutions during the textual encoding that target semantic disambiguation during encoding, rather than relying on downstream interventions such as Chefer et al.
+(2023a).
+Feng et al.
+(2023) offer one such solution, encoding items separately according to structured representations to improve attribute binding.
+While this represents a promising first step, future work must account for more complex interactions and employ more sophisticated, semantically-informed methods to achieve this separation.
+Figure 7: Verifying textual semantic leakage by replacing the contextually leaked concept representation.
+(1) Regular generation produces an image showing a crosswalk to the right of a bus station.
+(2) Generation from the prompt “standing zebra”, without any context, results in the correct interpretation of the zebra as an animal.
+(3) Generation using the original prompt, but with the leaked concept “standing zebra” replaced by it’s uncontextualized representation, yields a correct image.
+5 Related Work
+Interpretability in T2I models.
+Recent work has explored how T2I models encode and align concepts.
+Chefer et al.
+(2023b) studied CLIP’s latent space, while Toker et al.
+(2024) interpreted representations across the textual encoding process using the diffusion model as a lens.
+Others examined text-to-image alignment via attention maps (Tang et al., 2023).
+Another direction is to apply SAEs (Cunningham et al., 2023) to interpret intermediate representations in T2I models (Cywiński and Deja, 2025).
+We on the other hand, focus on token-level information at the encoder’s final layer, which directly conditions generation process.
+Token representation and flow in LLMs.
+Studies have shown that token information in LLMs is not uniformly distributed.
+Kaplan et al.
+(2025) found that subwords fuse into word-level meaning, while Feucht et al.
+(2024) documented how later tokens can erase earlier ones.
+Attention-based flow (Vig and Belinkov, 2019; Clark et al., 2019) can be misleading (Pruthi et al., 2020; Jain and Wallace, 2019), prompting alternatives like Attention Rollout (Abnar and Zuidema, 2020).
+Methods like Patchscopes (Ghandeharioun et al., 2024) and logit lens (nostalgebraist, 2020) decode internal representations, but do not test if downstream components use them.
+Probing and causal methods.
+Probing methods (Adi et al., 2016; Liu et al., 2019; Zhang and Bowman, 2018; Brunner et al., 2019) reveal information presence but are not fully reliable, as probes can exploit spurious correlations (Belinkov, 2022).
+We instead use causal interventions to test whether token information is actually used during image generation.
+Challenges in T2I models.
+Semantic leakage—when context distorts word meaning—has been observed in both LLMs (Gonen et al., 2024) and T2I (Rassin et al., 2022; Dahary et al., 2024).
+Another common issue is neglect, where key items are omitted (Chefer et al., 2023a; Chang et al., 2024b).
+While prior work focuses on solutions during the diffusion process, we show that some of these failures often originate in the text encoder.
+6 Discussion
+Improving generation via token-level intervention.
+Our results suggest that generation can be improved by removing non-representative tokens, which enhances prompt-image alignment and reduces noise.
+This intervention can be automated via a high-precision probe (see Section˜A.5), or integrated into model design, as in MrT5 (Kallini et al., 2025), which omits uninformative tokens in byte-level encoders (Xue et al., 2022).
+Similarly, attention-based techniques (Chefer et al., 2023a) may benefit from prioritizing representative tokens over full lexical spans.
+Developing textually challenging benchmarks.
+Current T2I datasets focus primarily on visual or spatial complexity (Huang et al., 2023a; Ghosh et al., 2023; Saharia et al., 2022; Yu et al., 2022b).
+Yet our findings show that even slight linguistic ambiguity—particularly with polysemous or compositional phrases—can cause encoding failures.
+This highlights the need for evaluation benchmarks that probe textual difficulty more directly, which may in turn drive improvements in encoder design.
+Generalization across T2I architectures.
+We repeat our main experiments on SDXL-Turbo and Sana (see A.3.2 and A.3.3), which use unidirectional text encoders, and on FLUX-dev (A.3.1), a compute-intensive variant of FLUX.
+We observe similar phenomena across models, indicating our findings are not specific to a particular T2I design.
+7 Conclusion
+Our investigation of text encoding in T2I models revealed two key properties: (1) lexical meaning tends to concentrate in one or two representative tokens rather than distributing evenly, and (2) contextual information can leak between lexical items, resulting in misinterpretation.
+Our analysis further indicates focused evaluation of the text encoder and token-level interventions are promising avenues towards more semantically aligned T2I generation.
+Limitations
+Evaluating token-level representations remains challenging.
+While we rely on strong vision-language models as judges and validate key findings through human evaluation, these are still approximations of true semantic alignment.
+Our prompt set focuses on object-centric, syntactically simple cases, which may limit generalization to prompts involving misspellings, rare words, or abstract concepts.
+Further work is needed to explore how information flow behaves under more linguistically complex conditions.
+Appendix A Appendix
+A.1 Technical Details
+A.1.1 Lexical Item Classification
+We define a lexical item as either a single word or a compound expression of multiple words that, in context, conveys a unified semantic meaning.
+A compound expression is treated as a single item when its words form a fixed lexical unit with cohesive semantics rather than merely exhibiting a modifier–head relationship.
+For example, while “broken mirror” describes a mirror’s state, expressions like “hot air balloon” or “teddy bear” denote entities with distinct identities.
+Similarly, although phrases such as “identical twins” or “baseball bat” might be interpreted as separate concepts, conventional usage supports their treatment as unified entities.
+We employ the reasoning model 03-mini-high as a classifier to tag multi-word lexical items in both the target prompts and the dataset.
+The model returns a list of identified multi-word expressions, while the remaining untagged words are treated as individual lexical items.
+A.1.2 Redundant Token Classification
+We propose a probing classifier to predict whether a token is redundant (i.e., non-representative of its lexical item) using solely its encoded representation, without the need to generate an image.
+For this purpose, we extracted the 6,966 tokens corresponding to 4,864 unique lexical items from our dataset.
+Each token was annotated with a binary label indicating whether it represents the lexical item by the VLM (see A.1.3 for more details).
+We split the data into training and validation sets using an 80-20 ratio.
+A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations.
+The results on the evaluation set are presented in Table˜3.
+The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
+Table 3: Performance of k-NN classifier ($k=5$) for predicting token redundancy
+Metric Accuracy Precision Recall F1-score
+Score 0.82 0.92 0.74 0.82
+A.1.3 Evaluation Visual Generations.
+We evaluate whether an image matches a textual description using Qwen2-VL-72B-Instruct (Wang et al., 2024). The following prompt is used:
+"In Yes, No and maybe.
+Does every image match one of those descriptions: (description string)?
+Answer Yes if all images match or relate to at least one description, Maybe if only some match, otherwise No."
+Here, the textual description can be either a single lexical item or a complete textual prompt.
+A.1.4 Evaluating Generated Textual Descriptions.
+We employ GPT-4o (OpenAI et al., 2024) to evaluate the textual interpretations produced by Patchscopes. We use the following prompt:
+“In Yes, No and Maybe. Does every image match the description: {Patchscopes_description} ? Answer Yes if all images match or relate to the description, Maybe if only some match, otherwise No.”
+A.1.5 Evaluating Relations Between Items.
+We enhance our leakage validation by distinguishing between cases where two lexical items exhibiting semantic leakage are perceptually bound together—for example, “old” and “man” in the prompt “a portrait of an old man”—and cases where they are not as “cone hat” and “eating” in the prompt “A person searing a cone hat is eating” (see Fig.˜6).
+To achieve this, we use a large language model (LLM) as a judge.
+Specifically, we use GPT-4o and employ the following prompt:
+“In Yes or No: in this prompt: {input_prompt}, are {item_1} and {item_2} perceptually bound together?”
+We then filter out all cases where the lexical items are perceptually bound together and find that only 6.5% instances exhibit unintentional leakage.
+A.1.6 Intended Item Evaluation
+Foe each lexical-item, we manualy create two interpretation - one in the intended interpretation in the prompt, and another is a possible wrong interpretation of the word in other contexts.
+For example, given the prompt “A standing zebra to the right of a city bus station”, the corrent interpretations would be “the animal zebra”, while the incorrect interpretations would be “A zebra crossing”.
+We then ask a VLM to evaluate the generated image, and asses if the first or the latter interpretations exists in the images.
+To evaluate the model’s capacity for contextual disambiguation, we manually define two interpretations for each lexical item in each prompt.
+The first is the intended semantic meaning derived from the prompt’s context, and the second is a alternative interpretation, that is wrong in this context.
+For instance, in the prompt “A standing zebra to the right of a city bus station”, the intended meaning of “zebra is the animal”, whereas the wrong interpretation is a “zebra crossing”.
+Subsequently, a VLM analyzes the generated image to determine if it depicts the intended interpretation or the wrong interpretation as we define it.
+A.1.7 Resources
+Our computational experiments involved inference with four distinct text-to-image models: Flux-dev, Flux-schnell, sdxl-turbo, and Sana.
+The parameter sizes for these models are approximately 12 billion for Flux-dev and Flux-schnell, 3.1 billion for sdxl-turbo, and a range of 0.6 to 4.8 billion for the Sana models, with our experiments utilizing a 1.6 billion parameter version.
+The total computational budget for these experiments is estimated to be approximately 480 GPU hours.
+The computing infrastructure consisted of a cluster of eight NVIDIA A100 GPUs.
+This configuration provided the necessary computational power for the large number of inference tasks performed.
+A.1.8 Use of AI Assistants.
+We utilized AI assistants to support this research.
+For coding the experiments, we used Microsoft’s Copilot and Anthropic’s Claude 3; all generated code was manually reviewed and validated by us to ensure it aligned with our requirements.
+For the paper, Google’s Gemini models (Pro and Flash) were used to improve the writing and clarity.
+We have carefully reviewed all content to ensure it accurately reflects our intentions.
+A.2 Data
+DrawBench (Saharia et al., 2022): We include all categories except for “misspelling”, “rare words”, and “text”. Overall we extract 134 prompts from DrawBench.
+Parti Prompts (Yu et al., 2022a): We include all categories except “Style & Format”, “Writing & Symbols”, and “Arts”.
+Overall we extract 923 prompts from Parti Prompts.
+The dataset is released under apache-2.0 license.
+In total, we obtain 1,056 prompts.
+Extended dataset of leakage prompts.
+Our augmentation process incorporates two components.
+First, we generate variations of existing prompts from Rassin et al.
+(2022) (e.g., modifying ’a gentleman with a bow in the forest’ to ’a man wearing a bow in the jungle’).
+Second, we introduce novel prompts with potential semantic leakage.
+For these prompts, we applied a one-lexical item change test by generating an image from a similar prompt that substitutes the affected or leaked item with an alternative term (e.g., replacing ’bishops’ with ’cardinals’ or ’checkers’ in ’chess in ’2 bishops playing chess’).
+This test ensures that minimal lexical modifications do not alter the intended semantic meaning while producing a different image due to semantic leakage from another item in the prompt (see the first two columns in Fig.˜8 for few visual examples).
+Together, these methods enrich the dataset and provide a robust framework for analyzing semantic leakage.
+The full list of prompts is available in an anonymous Git repository.
+[^3] https://anonymous.4open.science/r/ TokenRole-3E19
+Figure 8: Examples from our semantic leakage method.
+Left: standard generation of leakage contained prompt.
+Second: generation using a one-lexical item change test as part of the dataset creation (a minimal substitution to verify that a slight lexical change yields a different image).
+Third: image from the contextual representation (misinterpreted item).
+Forth: image from the uncontextualized representation (correct interpretation).
+Right: final generation after patching the correct, uncontextualized representation into the prompt.
+We split the data into training and validation sets using an 80-20 ratio.
+A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations.
+The results on the evaluation set are presented in Table˜3.
+The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
+A.3 Additional models
+A.3.1 FLUX-Dev
+In addition to our primary experiments with FLUX, we repeated all analyses using the Flux-dev variant.
+The redundant versus representative token experiments yielded similar trends, with 55% of tokens identified as representative and 45% as non-representative—values closely matching those observed with FLUX.
+Likewise, our inter-item flow experiments confirmed that information flow occurred in 11% of cases (and 3.1% miss intended leakage), reinforcing the overall patterns reported in the main text.
+Notably, while the aggregate trends are consistent across models, the specific lexical items resolved can differ between FLUX-schnell and Flux-dev, indicating potential a slightly different inner-lexicon (Kaplan et al., 2025).
+These findings underscore the robustness of our approach while highlighting model-dependent nuances in token representation and information flow dynamics.
+A.3.2 SDXL-Turbo
+Figure 9: Images generated from individual subtokens in SDXL-Turbo.
+We find that, in many cases, the representation of an item is not clearly reflected in any of its subtokens—for example, in the case of the token “skateboard.” Another interesting observation is that the last token of a lexical item often carries its representation, as seen in the “square” token of “times square.” We also observe that the EOT token incorporates information from the full prompt.
+Our analysis reveals that SDXL-Turbo, which uses the CLIP text encoder, behaves markedly differently from FLUX, which relies on the encoder in the encoder-decoder T5-XXL.
+In SDXL-Turbo, the text encoder is a causal language model, meaning each token’s encoding is influenced only by its preceding tokens during the encoding process.
+Figure 10: Token-level image generation using Sana on the prompt “a black baseball hat with a flame decal on it”.
+Representative information often resides in the last token of each item (e.g., “hat”, “decal”), consistent with Sana’s causal encoding.
+Contextual influence is one-directional, with earlier tokens shaping later ones.
+We repeated our in-item representation experiments using SDXL-Turbo.
+Our first observation is that most generated images are either abstract or unrelated to the intended lexical items.
+According to our analysis, 55% of lexical items in SDXL-Turbo lack any representative token (compared to just 11% in FLUX).
+Moreover, when a representative token is present in CLIP, it is typically the final token of the lexical item (see Fig.˜9).
+This is aligned with the unidirectional encoding of the model.
+Another phenomenon we observe—consistent with CLIP’s training objective—is the unusually dominant role of the end-of-sequence (EOS) token.
+Images generated from the EOS token often encapsulate nearly the full semantic content of the prompt.
+In our evaluation, 62% of EOS-generated images matched the prompt ( compared to 73% when using the full prompt).
+We believe this also causes our intervention method to be less effective, since when we interpret a single token, we patch all other tokens, including the EOT token—which usually contains a lot of information—with tokens derived from an empty prompt (see Fig.˜9).
+A.3.3 Sana (Gemma-based encoder)
+We also evaluate our methodology on the Sana model (Xie et al., 2024b), which uses a Gemma-based autoregressive language model as the text encoder.
+These results help validate the generality of our findings across architectures with differing encoding strategies.
+For in-item information flow, we observe that Sana produces fewer multi-token lexical items due to its larger vocabulary relative to T5 and CLIP.
+In cases where items do consist of multiple tokens, we find that only the last token typically acts as a representative token—a behavior aligned with the unidirectional nature of autoregressive encoders.
+In our cross-item information flow analysis, we find that 17.45% of item pairs in Sana exhibit contextual information flow, compared to 10.45% in FLUX.
+This increase likely stems from the one-sided (forward-only) nature of Sana’s encoding.
+For example, in the prompt “a black baseball hat with a flame decal on it”, we find contextual influence such as “black baseball” affecting “hat” and “flame” affecting “decal”, but not the reverse.
+See Fig.˜10 for illustrations.
+Overall, while Sana shows a slightly higher rate of contextual influence than FLUX, it preserves many of the key structural insights found in our primary analysis—particularly the sparsity and location of representative tokens.
+Unlike CLIP, which encodes substantial information in the EOS token, Sana lacks such artifacts, suggesting that EOS-related effects are not fundamental to autoregressive models more generally.
+A.4 Inter-Item Information Flow Framework
+To assess whether one lexical item encodes information about others in the same prompt, we conduct the following experiment.
+Given a prompt, we isolate each lexical item one at a time.
+First, we encode the full prompt using the text encoder.
+Then, for a given lexical item, we apply our patching method (Section˜2.1) to mask the representations of all other tokens—leaving only the contextualized representation of the target item intact.
+Formally, for a lexical item with token indices $S\subset\{1,\ldots,N\}$, we construct a patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$ in which only the tokens in $S$ retain their original contextualized representations.
+We then generate an image from this modified sequence, capturing what information is encoded in the selected item’s contextualized form.
+We repeat this process for the same item in an uncontextualized setting: we encode and generate an image using only that lexical item in isolation, without the rest of the prompt.
+This allows us to distinguish between information inherently present in the item’s encoding and information introduced by context.
+To measure influence between items, we use Qwen2-VL to check whether a second item $y$ appears in the image generated from a first item $x$, both in the contextualized and uncontextualized versions.
+Influence is considered True if $y$ appears in the image generated from contextualized $x$, but not from uncontextualized $x$.
+This comparison ensures that the observed presence of $y$ is attributable to contextual information flow during encoding, rather than coincidental co-occurrence or inherent correlation.
+This setup enables us to detect breaches of lexical segmentation—i.e., cases where one lexical item encodes visual information belonging to another—quantifying interdependence across items within the prompt.
+See Fig.˜11 for an illustration of the procedure.
 Figure 11: Evaluating inter-item information flow: Our proposed framework to interpret the information flow between lexical items in the prompt.
@@
 If this is the case, we conclude that there was information flow from “table” to “pool”.
-5 Inter-item information flow
-User prompts are typically composed of multiple lexical items.
-Here we investigate the influence of lexical items on one another.
-For example, in the sentence “a pool next to a table”, is the representation of the item “pool” affected by the presence of the item “table”?
-More generally, given a prompt, made up of several lexical items, we aim to measure whether the presence in the prompt of one item affects the representation of another.
-To this end, we conduct the following experiment: given a prompt, we isolate each lexical item one at a time.
-First, we encode the entire prompt.
-Next, for each item, we patch the representations of all other tokens using our method Section 3.1, and generate an image based on the modified sequence.
-Specifically, we define a subset $S\subset\{1,\ldots,N\}$ containing the token indices of the selected lexical item, construct the patched sequence $\tilde{t}_{1},\ldots,\tilde{t}_{N}$, and use it to guide the diffusion model.
-This process produces one image per lexical item in the prompt—each encoded in the full context of the prompt but with all other token representations masked.
-These images allow us to assess what information is encoded in each item’s representation.
-We then generate an image for the uncontextualized version of each lexical item—encoding only that item in isolation, without the rest of the prompt.
-To assess whether a lexical item $y$ influences another item $x$, we use Qwen2-VL to check if $y$ is present in the image generated from $x$, both with and without context.
-We then compare the Qwen2-VL answers between the contextualized and uncontextualized versions.
-Refer to A.1 for more details regarding Qwen2-VL and the prompts we use.
-An influence is considered True if $x$ did not exhibit $y$ in the uncontextualized representation but did in the contextualized one.
-Confirming that $y$ does not appear in the isolated representation of $x$ is crucial, as we want to ensure that the presence of $y$ in $x$ is due to information flow that occurs during textual encoding, and not due to some underlying correlation between the items.
-This approach allows us to quantify the degree to which lexical items affect one another within a prompt.
-See 4 for a demonstration of this framework.
-Figure 5: Examples of information flow between items.
-Top: Images generated from a lexical item encoded alongside another item that alters its representation.
-Bottom: Images generated from the uncontextualized representation of the same lexical item.
-The first three images (from the left) demonstrate correct information flow, while the last image (far right) demonstrates incorrect information flow.
-Our results show that in 89% of the cases, lexical items are not influenced by other items in the prompt.
-In the remaining 11%, one lexical item incorporates information about another item that appears in the same prompt; see 5 for examples.
-Interestingly, we find that when such influence occurs, it often involves items that frequently co-occur in natural language, and are therefore likely to appear together in the training data.
-For example, the item “pool” often shifts its representation from a swimming pool to a pool table when “table” is present in the context.
-This phenomenon occurs even when “table” is not used as a pool table, as in the sentence “A pool next to a table” (see 4).
-To measure the extent of information flow between items that should not influence each other (as in the “pool” and “table” example), we divide the identified cases of flow between items into two categories.
-The first includes cases where the items are semantically linked, such as “black” and “bear” in the prompt “a black bear”.
-The second includes cases where the items are not semantically linked, such as “black” and “bear” in the sentence “a green bear by the black tree”.
-To achieve this, we use an LLM as a judge (details in A.1).
-Our analysis reveals that 31% out of the influenced instances exhibit unintentional information flow.
-Application.
-Based on our findings, we present a method for mitigating semantic leakage.
-Our approach first encodes the prompt as usual, then separately encodes the item suspected of leakage.
-Finally, we replace the representation of the contextualized, leaked item in the original prompt’s encoding with its uncontextualized representation, and condition the diffusion process on this mixed representation (see 6).
-To evaluate the effectiveness of our method, we adapt prompts from Rassin et al.
-(2022) as these prompts often induce semantic leakage.
-For example, given the prompt “a bat is flying over a baseball stadium”, models tend to generate a flying baseball bat rather than the animal, due to the strong association between “bat” and “baseball stadium”.
-Since the dataset contains only about 30 examples, we use GPT-4o to expand it to 110 prompts.
-See A.2 for details on the augmentation process, including the full list of prompts.
-Out of the 110 prompts, 104 exhibit semantic leakage (as determined by manual human evaluation), where the image contains a concept not intended by the prompt.
-We then select the lexical items in each prompt that we suspect of having leaked, and apply our method on them.
-Remarkably, we achieve an 85% improvement on this set.
-See 8 in A.3 for additional qualitative examples.
-Figure 6: Removing semantic leakage by replacing the contextually leaked concept representation.
-(1) Regular generation produces an image showing a crosswalk to the right of a bus station.
-(2) Generation from the prompt “standing zebra”, without any context, results in the correct interpretation of the zebra as an animal.
-(3) Generation using the original prompt, but with the leaked concept “standing zebra” replaced by it’s uncontextualized representation, yields a correct image.
-6 Conclusions
-In this study, we examined the information flow within lexical items and between different lexical items during the textual encoding of a prompt in T2I models.
-We introduced an intervention to evaluate how information is distributed across various token representations.
-For in-item information flow, our analysis revealed that information within a concept is not uniformly distributed; instead, it is primarily concentrated in a few specific tokens.
-For inter-item information flow, we found that information typically does not flow between different items.
-When it does, the flow is often unintended and may result in semantic leakage.
-Our experiments suggest that semantic leakage originates within the text encoder itself, challenging the common assumption that each token primarily captures its intended meaning.
-We showed that semantic leakage phenomenon can be largely mitigated by simply encoding the leaked item separately and patching it into the prompt representation before the diffusion.
-While prior work has focused on modifying cross-attention to address challenges in text-to-image models, our results suggest that these methods should also account for information flow between tokens that occurs earlier, during the textual encoding.
-7 Acknowledgments
-This research was supported in part by the Israel Science Foundation (grant No.
-448/20), NSF-BSF grant 2020793, an Azrieli Foundation Early Career Faculty Fellowship, an AI Alignment grant from Open Philanthropy, the European Union (ERC, Control-LM, 101165402), and by an Academic Gift by NVIDIA.
-Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
-Neither the European Union nor the granting authority can be held responsible for them.
-Appendix A Appendix
-A.1 Technical details
-Lexical item classification
-We define a lexical item as either a single word or a compound expression of multiple words that, in context, conveys a unified semantic meaning.
-A compound expression is treated as a single item when its words form a fixed lexical unit with cohesive semantics rather than merely exhibiting a modifier–head relationship.
-For example, while “broken mirror” describes a mirror’s state, expressions like “hot air balloon” or “teddy bear” denote entities with distinct identities.
-Similarly, although phrases such as “identical twins” or “baseball bat” might be interpreted as separate concepts, conventional usage supports their treatment as unified entities.
-We employ the reasoning model 03-mini-high as a classifier to tag multi-word lexical items in both the target prompts and the dataset.
-The model returns a list of identified multi-word expressions, while the remaining untagged words are treated as individual lexical items.
-Redundant token classification
-We propose a probing classifier to predict whether a token is redundant (i.e., non-representative of its lexical item) using solely its encoded representation, without the need to generate an image.
-For this purpose, we extracted the 6,966 tokens corresponding to 4,864 unique lexical items from our dataset.
-Each token was annotated with a binary label indicating whether it represents the lexical item by the VLM (see A.1 for more details).
-We split the data into training and validation sets using an 80-20 ratio.
-A $k$-nearest neighbors (k-NN) classifier with $k=5$ and Euclidean distance as the similarity measure was then applied to predict token redundancy directly from the encoded representations.
-The results on the evaluation set are presented in 1.
-The high precision indicates that, in practical settings, one can remove tokens predicted as redundant with a high degree of certainty that representative tokens will not be inadvertently discarded.
-Table 1: Performance of k-NN classifier ($k=5$) for predicting token redundancy
-Metric Accuracy Precision Recall F1-score
-Score 0.82 0.92 0.74 0.82
-Evaluation visual generations.
-We evaluate whether an image matches a textual description using Qwen2-VL-72B-Instruct (Wang et al., 2024). The following prompt is used:
-”In Yes, No and maybe.
-Does every image match one of those descriptions: (description string)?
-Answer Yes if all images match or relate to at least one description, Maybe if only some match, otherwise No.”
-Here, the textual description can be either a single lexical item or a complete textual prompt.
-Evaluating generated textual descriptions.
-We employ gpt4o (OpenAI et al., 2024) to evaluate the textual interpretations produced by Patchscopes. We use the following prompt:
-“In Yes, No and Maybe. Does every image match the description: {Patchscopes_description} ? Answer Yes if all images match or relate to the description, Maybe if only some match, otherwise No.”
-Classifying the relations between items.
-We enhance our leakage validation by distinguishing between cases where two lexical items exhibiting semantic leakage are perceptually bound together—for example, “old” and “man” in the prompt “a portrait of an old man”—and cases where they are not as “cone hat” and “eating” in the prompt “A person searing a cone hat is eating” (see 5).
-To achieve this, we use a large language model (LLM) as a judge.
-Specifically, we employ the following prompt:
-“In Yes or No: in this prompt: {input_prompt}, are {item_1} and {item_2} perceptually bound together?”
-We then filter out all cases where the lexical items are perceptually bound together and find that only 6.5% instances exhibit unintentional leakage.
-A.2 Data
-DrawBench (Saharia et al., 2022): We include all categories except for “misspelling”, “rare words”, and “text”.
-Overall we end up with 134 prompts from DrawBench.
-Parti Prompts (Yu et al., 2022a): We include all categories except “Style & Format”, “Writing & Symbols”, and “Arts”.
-Overall we end up with 923 prompts from Parti Prompts.
-In total, we end up with 1,056 prompts.
-Extended dataset of leakage prompts
-Our augmentation process incorporates two components.
-First, we generate variations of existing prompts from Rassin et al.
-(2022) (e.g., modifying ’a gentleman with a bow in the forest’ to ’a man wearing a bow in the jungle’).
-Second, we introduce novel prompts with potential semantic leakage.
-For these prompts, we applied a one-lexical item change test by generating an image from a similar prompt that substitutes the affected or leaked item with an alternative term (e.g., replacing ’bishops’ with ’cardinals’ or ’checkers’ in ’chess in ’2 bishops playing chess’).
-This test ensures that minimal lexical modifications do not alter the intended semantic meaning while producing a different image due to semantic leakage from another item in the prompt (see the first two columns in 8 for few visual examples) .
-Together, these methods enrich the dataset and provide a robust framework for analyzing semantic leakage.
-The full list of prompts is available in the following anonymous Git repository: https://anonymous.4open.science/r/TokenRole-3E19.
-A.3 Additional results
-Figure 7: Distribution of representative vs. non-representative tokens across different lexical item lengths.
-Method Success Rate
-Regular Generation 84.1%
-After Removing Redundant Tokens 87.6%
-Table 2: Image-text alignment success rate (evaluated with Qwen-VL): comparison between standard generation and generation with redundant tokens removed
-Figure 8: Examples from our semantic leakage method.
-Left: standard generation of leakage contained prompt.
-Second: generation using a one-lexical item change test as part of the dataset creation (a minimal substitution to verify that a slight lexical change yields a different image).
-Third: image from the contextual representation (misinterpreted item).
-Forth: image from the uncontextualized representation (correct interpretation).
-Right: final generation after patching the correct, uncontextualized representation into the prompt.
-A.4 Catastrophic negligence
-Catastrophic negligence refers to cases where a concept or entity is completely absent from the generated image.
-Chefer et al.
-(2023a) attempted to address this issue by increasing attention to the neglected item, implicitly assuming that the model understands the concept.
-We find that in 7% of the cases, the item is not generated from any of its constituent tokens and is absent from the full prompt’s image.
-To better understand this phenomenon, we leverage Patchscopes (Ghandeharioun et al., 2024), a tool that interprets intermediate token representations by generating textual explanations using the decoder in encoder-decoder architectures.
-Using Patchscopes, we observe that in approximately 67% of these cases the tokens are explained coherently, indicating that the item is adequately represented in the text encoder.
-For example, when we patch the encoded representation of the lexical item tuba into the prompt “describe X”, Patchscopes describes it as “a musical instrument played by blowing into the mouthpiece.” However, when we try to generate an image from its representation, the model fails (see 9), suggesting that the diffusion model may have encountered too few examples of the intended entity during training.
-Additionally, we identify two abstract categories that consistently result in failure in image generation, while Patchscopes accurately describes the concept: negation and numbers.
-For instance, the lexical item “without” failed to generate the an image without the item in 100% of 100 observed instances.
-See  9.
-Figure 9: Cases where intervention-based interpretations fail, while Patchscopes accurately describe token meaning.
-Each column corresponds to a different prompt.
-The first row shows images generated from the contextual bolded token.
-The second row presents the same token explained by Patchscopes, and the third row shows the final image generated from the full prompt.
-Left: The image generated from the token “tub” (from “tuba”) depicts a lamp, whereas Patchscopes correctly identifies it as a musical instrument.
-Middle: The concept “without” fails to yield an image lacking a refrigerator, although Patchscopes correctly describes its semantics.
-Right: The concept “7” leads to an incorrect number of entities in the image, while Patchscopes correctly identifies the quantity.
-A.5 Additional models
-A.5.1 FLUX-Dev
-In addition to our primary experiments with FLUX, we repeated all analyses using the Flux-dev variant.
-The redundant versus representative token experiments yielded similar trends, with 55% of tokens identified as representative and 45% as non-representative—values closely matching those observed with FLUX.
-Likewise, our inter-item flow experiments confirmed that information flow occurred in 11% of cases (and 3.1% miss intended leakage), reinforcing the overall patterns reported in the main text.
-Notably, while the aggregate trends are consistent across models, the specific lexical items resolved can differ between FLUX-schnell and Flux-dev, indicating potential a slightly different inner-lexicon.
-These findings underscore the robustness of our approach while highlighting model-dependent nuances in token representation and information flow dynamics.
-A.5.2 SDXL-Turbo
-Figure 10: Images generated from individual subtokens in SDXL-Turbo.
-We find that, in many cases, the representation of an item is not clearly reflected in any of its subtokens—for example, in the case of the token “skateboard.” Another interesting observation is that the last token of a lexical item often carries its representation, as seen in the “square” token of “times square.” We also observe that the EOT token incorporates information from the full prompt.
-Our analysis reveals that SDXL-Turbo, which uses the CLIP text encoder, behaves markedly differently from FLUX, which relies on the encoder in the encoder-decoder T5-XXL.
-In SDXL-Turbo, the text encoder is a causal language model, meaning each token’s encoding is influenced only by its preceding tokens during the encoding process.
-We repeated our in-item information flow experiments using SDXL-Turbo.
-Our first observation is that most generated images are either abstract or unrelated to the intended lexical items.
-According to our analysis, 55% of lexical items in SDXL-Turbo lack any representative token (compared to just 11% in FLUX).
-Moreover, when a representative token is present in CLIP, it is typically the final token of the lexical item (see 10).
-This is aligned with the unidirectional encoding of the model.
-Another phenomenon we observe—consistent with CLIP’s training objective—is the unusually dominant role of the end-of-sequence (EOS) token.
-Images generated from the EOS token often encapsulate nearly the full semantic content of the prompt.
-In our evaluation, 62% of EOS-generated images matched the prompt ( compared to 73% when using the full prompt).
-We believe this also causes our intervention method to be less effective, since when we interpret a single token, we patch all other tokens, including the EOT token—which usually contains a lot of information—with tokens derived from an empty prompt (see 10).
+A.5 Predicting Redundant Tokens Without Image Generation
+Identifying which tokens are representative can improve image generation, as demonstrated in Fig.˜4 (right).
+Moreover, many existing T2I methods operate at the token level, typically treating all of a lexical item’s tokens equally (Chefer et al., 2023a; Dahary et al., 2024).
+However, our findings suggest that only a subset of these tokens contribute meaningfully to the representation.
+Yet, identifying representative tokens via image generation is computationally expensive. To address this, we explore alternatives for estimating token importance without image-level interventions.
+First, we compute the character-level edit distance between a lexical item (e.g., “giraffe”) and its constituent tokens (e.g., “gir”, “a”, “ffe”).
+We observe a negative correlation between edit distance and token representativeness (Pearson correlation of 0.44-0.44), suggesting that tokens more orthographically distant from the original item are less likely to be representative.
+While informative, this signal alone is insufficient for accurate prediction.
+Instead, we train a $k$-Nearest Neighbors (k=5k{=}5) classifier using Euclidean distance over the token representations to predict redundancy.
+Using 6,966 token instances (spanning 4,864 unique lexical items), each labeled for whether the token is representative according to a VLM, our classifier achieves 92% precision in identifying non-representative tokens.
+This result highlights the feasibility of lightweight approaches for filtering redundant tokens without resorting to costly generation procedures.
+See Section˜A.1.2 for classifier details.